OpenStudy (deltanu):

Problem set 7 3C-4 asks us to find the center of mass of a sector of given angle. Can anyone explain why it isn't possible to find the 'polar' centre of mass by simply integrating r*rdrdθ? I get that it won't come out dependent on θ in this case, but I don't understand why we can't find COM in polar coordinates as we can in rectangular... I'm fond of intuitive geometrical explanations but I can't think of one!

OpenStudy (phi):

what makes you think you can't?

OpenStudy (deltanu):

Because when we naively try to use the equation \[r _{cm}=\int\limits_{0}^{2\alpha}\int\limits_{0}^{a}\delta\ r*rdrd \theta \] we get a nonsense answer of 2a/3. I'm just wondering why it is we can't do this in the same way that we found xcm...

OpenStudy (baru):

the formula for \(x_{cm}\) is derived by balancing moments generated by masses about an axis parallel to the y axis

OpenStudy (baru):



OpenStudy (baru):



OpenStudy (baru):

we want to find \(x_cm\) such that \(m_1r_1+m_2r_2=0\) \[m_1(x_{cm}-x_1) +m_2(x_{cm}-x_2)=0\]

OpenStudy (baru):

\[(m_1+m_2)x_{cm}=m_1x_1+m_2x_2\] \[x_{cm}=\frac{m_1x_1+m_2x_2}{(m_1+m_2)}\]

OpenStudy (baru):

its the same derivation which can be extended to continuously distributed mass... if you replace x with r or theta, it just dosent make sense

OpenStudy (baru):

because the idea of balancing moments is lost

OpenStudy (deltanu):

Great explanation. That really clears things up! Definitely changed the way I think about centre of mass. The concept of moments is key to the whole definition... Thanks for the help

OpenStudy (baru):

no prob :)

OpenStudy (baru):

just to clarify: you can still switch to polar and find \(x_{cm}\) (not \(r_{cm}\)) \[x_{cm}=\int\limits\limits\int\limits\limits_{R}\delta\ rcos \theta*rdrd \theta\]